A random sample of size 25 was taken from a normal population with σ 2 = 6. A confidence interval for the mean was given as (5.37, 7.37). What is the confidence coefficient associate with this interval?
Answer #1

Margin of Error = (Upper-lower)/2 = (7.37-5.37)/2 = 1

Ans : 0.958 = 95.8 %